the given conditions to write an equation for the line in point-slope form.
{{slope=6/7, passing through (8,7)}} is the answer y+7=6/7 (x+8)

Answer by nerdybill(7384)

the given conditions to write an equation for the line in point-slope form.
{{slope=6/7, passing through (8,7)}} is the answer y+7=6/7 (x+8)
.
"point-slope" form is:
y - y1 = m(x - x1)
plugging in given values:
y - 7 = 6/7(x - 8)